OpenStudy (anonymous):

What is sin30 = x / 45

OpenStudy (anonymous):

21

OpenStudy (anonymous):

jk im sorry lol i had to hahaha im laughing so hard

OpenStudy (anonymous):

Oh.. aight lol

OpenStudy (anonymous):

@Skielerlucas04 lolll

OpenStudy (anonymous):

lol :D

Vocaloid (vocaloid):

sin(30) = ?

Vocaloid (vocaloid):

(use a calculator or the unit circle)

OpenStudy (anonymous):

[\sqrt{3} over 2

OpenStudy (anonymous):

or 0.8660254038

OpenStudy (anonymous):

Hm, I'll just give the full question lol A 45 foot long escalator rises from the first floor to the second floor of a shopping mall. The escalator makes a 30° angle with the horizontal. How high above the first floor is the second floor? A. 26ft B. 78ft C. 22.5ft D. 45ft

Vocaloid (vocaloid):

solve for x sin30 = x / 45 first, multiply both sides by 45

Vocaloid (vocaloid):

whatever you get for x will be the final answer

OpenStudy (anonymous):

that a 30, 60, 90 triangle

OpenStudy (anonymous):

I got -44.46 = x

OpenStudy (anonymous):

so far u are right

OpenStudy (anonymous):

Oh fr? okay thanks !

OpenStudy (anonymous):

im pretty sure its 26

OpenStudy (anonymous):

Oh, how did you get that?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so confused rn, lol

OpenStudy (anonymous):

me too

OpenStudy (anonymous):

ya know how 30, 60, 90 triangles the sides are 3,4,5

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

so i divided 45 by 5 and got 9

OpenStudy (anonymous):

do u see why?

OpenStudy (anonymous):

I see how you got 45 / 5 = 9 but then how does it turn to 26ft

OpenStudy (anonymous):

cause u multiply 9x3=27

OpenStudy (anonymous):

but obviously its not gonna be exact so i rounded it down to 26

OpenStudy (anonymous):

ya gotta know hot to cheat math sometimes lol

OpenStudy (anonymous):

Lol, aight thanks!

OpenStudy (anonymous):

u got it?

OpenStudy (anonymous):

oh wait, so when you did 9 x 3, wouldn't it have been 9 x 4 since its on the 60 degree angle side

OpenStudy (anonymous):

no because the side x is opposite the angle 30

OpenStudy (anonymous):

Ohhhhhh I see, aight I got it now, thanks bro

OpenStudy (anonymous):

tell me if its right

OpenStudy (anonymous):

wait so you don't know if its right? lmao

OpenStudy (anonymous):

lmao im 99% sure

OpenStudy (anonymous):

99.9% actually

OpenStudy (anonymous):

xD well I gotta do like 7 more questions to see if its right

OpenStudy (anonymous):

lol ok

OpenStudy (anonymous):

thats what i got

OpenStudy (anonymous):

So uhm, it wasn't correct.. lmao bruh